# UrbanKGent: A Unified Large Language Model Agent Framework for Urban Knowledge Graph Construction

### v1

@@ 5. Experiments
Pretrained Language Models KG-BERT - 0.24/3.15 - 0.23 - 0.19/4.12 - 0.24
@@
Zero-shot Reasoning Llama-2-7B 0.14/1.98 0.18/3.75 0.16 0.18 0.26/1.96 0.15/2.83 0.21 0.22
@@
In-context Learning Llama-2-7B (1-shot) 0.12/2.07 0.19/3.91 0.15 0.16 0.21/2.07 0.16/3.09 0.20 0.19
Llama-2-7B (3-shot) 0.18/2.15 0.21/3.96 0.19 0.18 0.25/2.44 0.18/3.27 0.23 0.20
Llama-2-13B (1-shot) 0.22/2.45 0.26/3.87 0.20 0.19 0.24/3.09 0.15/2.98 0.21 0.23
Llama-2-13B (3-shot) 0.26/3.52 0.31/3.28 0.23 0.24 0.28/2.65 0.21/2.53 0.25 0.26
GPT-3.5 (1-shot) 0.32/4.25 0.38/3.87 0.37 0.28 0.33/3.88 0.30/4.17 0.36 0.34
GPT-3.5 (3-shot) 0.41/4.65 0.42/4.08 0.42 0.31 0.36/4.24 0.36/4.23 0.39 0.36
@@
UrbanKGent Inference Llama-2-7B 0.27/3.05 0.26/4.12 0.28 0.24 0.27/2.87 0.24/3.54 0.26 0.29
@@
5.1. Experimental Settings
5.1.1. Dataset
In this work, two sequential tasks (i.e., RTE and KGC) of UrbanKGC are within an open-world setting (i.e., no predefined ontology) (Lu et al., 2023; Ye et al., 2022). Therefore, for the RTE task, every data record is an urban text without the triplet label. For the KGC task, every data record is a quadruple (i.e., head entity name, head entity geometry, tail entity name, tail entity geometry) without the geospatial relation label. We construct the RTE and KGC datasets of NYC and CHI by sampling uniformly from five raw data in Table 1, respectively. As shown in Table 2, we first construct two small datasets (i.e., NYC-Instruct and CHI-Instruct) for instruction fine-tuning and two middle datasets (i.e., NYC and CHI) to validate the performance of the constructed UrbanKGC agent. The remaining data works as the large-scale UrbanKGC dataset (i.e., NYC-Large and CHI-Large) in real-world scenarios shown in Table 7. The three types of datasets are non-overlapping to prevent data leakage. More dataset construction details are in Appendix A.
For each UrbanKGC dataset, we follow Section 4.5 to prompt constructed UrbanKGent to complete RTE and KGC tasks.
5.1.2. Baseline Methods
We provide a comprehensive comparison of our method with existing paradigms (e.g., zero-shot reasoning (Wang et al., 2023a), in-context learning (Wei et al., 2022) and vanilla finetuning (Yao et al., 2023)) on the UrbanKGC tasks.
Pretrained Language Model Baselines. For the RTE task, we utilize RelationPrompt (Chia et al., 2022), an end-to-end generation model for zero-shot RTE. For the KGC task, we fine-tune KG-BERT (Yao et al., 2019) and KG-T5 (Saxena et al., 2022) with the QA pairs constructed from the self-instruct dataset.
LLMs-based Zero-shot Reasoning Methods. We directly prompt the LLMs with basic task definitions to get the answer without training.
LLMs-based In-context Learning Methods. We sample a few-shot QA pairs as demonstrations from the self-instruct dataset as examples and get the answers from the LLMs without training.
Vanilla Fine-tuning Methods. We directly fine-tune the LLMs using the QA pairs constructed from the self-instruct dataset, and then prompt the LLMs with basic task definition without demonstrations.
UrbanKGent Inference Methods. We directly prompt the LLMs using the UrbanKGgent inference pipeline in Section 4.5.
The prompt templates of the above baseline methods are in Appendix B.
5.1.3. Implementation and Detail Settings
In our experiment, we select Llama-2, GPT-3.5 and GPT-4 as the backbone LLM $\mathcal{M}$. We follow the open-source Llama document [^8] https://ai.meta.com/llama/ to deploy Llama-2-7B, Llama-2-13B, and Llama-2-70B in our device. We implement two fine-tuning agent versions (e.g., Llama-2-7B and Llama-2-13B), and all experiments are conducted on eight NVIDIA A800 GPUs. For the GPT-3.5 and GPT-4, we adopt the gpt-3.5-turbo-16k-0613 API and gpt-4-0613 API from the GPT service [^9] https://gpt.hkust-gz.edu.cn/ provided by HKUST (GZ).
5.1.4. Evaluation Protocol
@@
Since the UrbanKGC tasks in this work follow an open-world setting where labels are not visible, the classical metric (e.g., F1 and Hits@10) is not applicable. In this work, we regard evaluation as the binary classification, i.e., if the extracted triplet in RTE task is correct and if the completed relation in KGC task is correct. We follow recent LLMs-based KGC works (Yao et al., 2023) to employ accuracy as an evaluation metric. To make a comprehensive evaluation of the experimental results, we employ the evaluation from two aspects:
Human Evaluation. We employ human annotators to evaluate the results on 200 random samples. For the relational triplet extraction task, we first manually annotate the triplet label for each sample. Then, we manually evaluate the correctness of each triplet (Lu et al., 2023) based on annotation and calculate the accuracy value. For the knowledge graph completion task, we follow (Yao et al., 2023) to manually label the response as correct or wrong, and calculate the accuracy. The detailed human evaluation process is in Appendix D.
@@
GPT Evaluation. Recently, many studies (Zheng et al., 2023; Wang et al., 2023a) adopt LLM-based evaluation for open-domain tasks and empirically demonstrate that GPT-4’s evaluation and human evaluation can be consistent (Cheng et al., 2023). In this work, we also use GPT-4 to evaluate the model performance on the full data to escape intensive labor. Specifically, given an UrbanKGC instruction and results, we prompt GPT-4 to return the confidence score and the justification (i.e., True/False), which will be further used to calculate the accuracy. In this work, the GPT-4’s evaluation has also been demonstrated to be consistent with the human evaluation. The detail is in Appendix D.
Table 4. Effect of different blocks.

### v2

@@ 5.1 Experimental Settings
End-to-end Models KG-BERT - 0.24/3.15 - 0.23 - 0.19/4.12 - 0.24
@@
PRGC 0.08/4.01 - 0.06 - 0.13/4.15 - 0.15 -
Zero-shot Reasoning Vicuna-7B 0.12/2.84 0.19/4.06 0.14 0.16 0.22/4.12 0.14/4.03 0.21 0.18
Alpaca-7B 0.11/3.75 0.17/3.87 0.15 0.17 0.23/3.96 0.16/4.15 0.20 0.16
Mistral-7B 0.14/4.12 0.21/4.11 0.17 0.18 0.21/3.75 0.15/3.76 0.19 0.19
Llama-2-7B 0.14/1.98 0.18/3.75 0.16 0.18 0.26/1.96 0.15/2.83 0.21 0.22
Llama-3-8B 0.15/4.02 0.15/4.02 0.20 0.21 0.24/3.75 0.15/4.08 0.22 0.22
@@
In-context Learning Llama-2-7B 0.18/2.15 0.21/3.96 0.19 0.18 0.25/2.44 0.18/3.27 0.23 0.20
Llama-3-8B 0.17/4.06 0.18/3.53 0.21 0.22 0.28/4.31 0.17/4.14 0.24 0.21
Llama-2-13B 0.26/3.52 0.31/3.28 0.23 0.24 0.28/2.65 0.21/2.53 0.25 0.26
GPT-3.5 0.41/4.65 0.42/4.08 0.42 0.31 0.36/4.24 0.36/4.23 0.39 0.36
@@
UrbanKGent Inference Vicuna-7B 0.24/3.07 0.24/3.95 0.29 0.23 0.27/4.12 0.22/3.95 0.23 0.25
Alpaca-7B 0.26/3.85 0.27/3.83 0.26 0.22 0.27/3.83 0.21/4.12 0.27 0.29
Mistral-7B 0.26/4.15 0.25/4.08 0.28 0.23 0.25/3.61 0.21/4.08 0.25 0.26
Llama-2-7B 0.27/3.05 0.26/4.12 0.28 0.24 0.27/2.87 0.24/3.54 0.26 0.29
Llama-3-8B 0.29/4.15 0.31/4.08 0.33 0.26 0.26/3.28 0.24/3.97 0.30 0.31
@@
UrbanKGent-8B 0.47/3.97 0.51/4.15 0.49 0.45 0.49/3.97 0.44/4.05 0.55 0.46
$\uparrow 4.44\%$ $\uparrow 6.25\%$ $\uparrow 4.26\%$ $\uparrow 7.14\%$ $\uparrow 6.52\%$ $\uparrow 7.32\%$ $\uparrow 5.77\%$ $\uparrow 6.98\%$
@@
Baseline Methods. We provide a comprehensive comparison of our method with existing paradigms: (1) End-to-end Models: For the zero-shot RTE task, we utilize the end-to-end generation model RelationPrompt [42] and PRGC [10]. For the KGC task, we fine-tune KG-BERT [43] and KG-T5 [44] with the QA pairs constructed from the self-instruct dataset. (2) LLMs-based Zero-shot Reasoning [45]: We directly prompt the LLMs with basic task definitions to get the answer without training. (3) LLMs-based In-context Learning [35]: We sample 3-shot QA pairs as demonstrations from the self-instruct dataset as examples and get the answers from the LLMs without training. (4) Vanilla Fine-tuning [11]: We directly fine-tune the LLMs using the QA pairs constructed from the self-instruct dataset, and then prompt the LLMs with basic task definition without demonstrations. (5) UrbanKGent Inference: We directly prompt the LLMs using the UrbanKGgent inference pipeline in Section 4.5. The prompt templates of the above baseline methods are in Appendix B.
Implementation and Detail Settings. In our experiment, we select Vicuna [46], Alpaca [47], Mistrial [48], Llama-2 [49], Llama-3 [50], GPT-3.5 [51] and GPT-4 [51] as the backbone LLM $\mathcal{M}$. All experiments are conducted on eight NVIDIA A800 GPUs. For the GPT-3.5 and GPT-4, we adopt the gpt-3.5-turbo-16k-0613 API and gpt-4-0613 API.
@@
Evaluation Protocol. Since UrbanKGC tasks in this work follow an open-world setting where labels are not visible, the classical metric (e.g., F1 and Hits@10) is not applicable. In this work, we regard evaluation as the binary classification, i.e., if the extracted triplet in RTE task is correct and if the completed relation in KGC task is correct. We follow recent LLMs-based KGC works [11] to employ accuracy as an evaluation metric. To make a comprehensive evaluation of the experimental results, we employ both of the human evaluation and GPT evaluation, which has been widely used in many LLM studies [52, 45]. For Human Evaluation, we employ human annotators to evaluate the results on
Table 4: Statistics comparison of constructed UrbanKGs in New York and Chicago between UrbanKGent and existing benchmark.
Dataset # Entity # Relation # Triplet Data Volume
NYC-Large 228,928 2,138 905,442 40,480
CHI-Large 95,813 1,336 563,290 28,607
NYC-UUKG 236,287 13 930,240 236,277
CHI-UUKG 140,602 13 564,400 140,577
200 random samples. As for the GPT Evaluation, we use GPT-4 to evaluate the model performance on the full data to escape intensive labor. In this work, the GPT-4’s evaluation has been demonstrated to be consistent with the human evaluation. Detail is in Appendix D.
5.2 Main Result
The performance results are reported in Table 3. As can be seen, the constructed agent outperforms all thirty-one baseline models on two UrbanKGC datasets. Specifically, the UrbanKGent-13B achieves (15.56%, 14.29%, 14.89%, and 11.90%) improvements compared with the state-of-the-art GPT-4 with the same inference pipeline on NYC. The improvements on CHI are (15.22%, 17.07%, 13.46%, and 13.95%), respectively. Moreover, the UrbanKGent-7/8B also achieve comparable performance compared with the GPT-4.
Meanwhile, we observe that the zero-shot LLMs perform poorly in the UrbanKGC tasks, even using GPT-4. In addition, although the demonstrations provided by In-context-learning can incorporate the UrbanKGC task information, the performance gain is limited. Besides, we find that fine-tuning LLMs can make obvious improvements in the overall performance. Through vanilla fine-tuning, the Llama-2-7/13B and Llama-3-8B could achieve comparable performance with GPT-3.5 under the ZSL settings.
Moreover, although the various LLM backbones using the UrbanKGent inference pipeline perform slightly worse than the vanilla fine-tuning method, they could obtain better performance compared
Figure 5: The model latency and cost of constructed UrbanKGent-13B and GPT-4 in UrbanKGC. We report the total inference time and cost of 1,000 RTE and KGC tasks.
with zero-shot reasoning and In-context learning paradigms. Such results demonstrate the benefit of knowledgeable instruction design and external tool innovation, but also indicate its performance bottleneck. As a deeper exploration, our work fills this gap through hybrid instruction fine-tuning, and the fine-tuned UrbanKGC agents, whether 7B, 8B or 13B, can achieve state-of-the-art performance in UrbanKGC tasks. We provide an in-depth analysis of the proposed UrbanKGent framework in Appdendix E.2.
5.3 Agent Application
We first derive UrbanKGent-13B for initial UrbanKGs acquisition in New York City and Chicago. After proper filtering and merging of the triplets, we obtain two large-scale UrbanKGs shown in Table 4. Compared with existing UrbanKG benchmark [8], we only use roughly one-fifth of the data for constructing the UrbanKGs with the same scale of triplets and entities, and even expanding the variety of relationships to a hundred times the original types. Moreover, we also provide efficiency analysis in Figure 5. As can be seen, UrbanKGent-13B achieves lower inference speed in latency and reduce the cost by roughly 20 times in both of RTE and KGC tasks. More details is in Appendix E.3.
6 Related work
Domain-Oriented Agent Construction. The concept of language agent [34] has become very popular recently, and a variety of LLM agents targeting different domains have been proposed. For example, Voyager [16] is constructed for automated game exploration, WebGPT [17] is an HTML agent for diverse document understanding tasks, LLMLight [53] constructs a language agent for transportation domain, K2 [13], GeoGalactica [19] and GeoLLM [12] propose to re-train language agent for geospatial semantic understanding. In addition, many recent works like Auto-GPT [54] and CAMEL [55] aim at proposing an autonomous agent framework for agent construction. Nevertheless, there is still no UrbanKGC agent construction framework for the urban computing domain.
LLMs for Knowledge Graph Construction. Recently, the advent of LLMs [56] invigorated the field of NLP. Many studies have begun to explore the potential of LLMs in the domain of KG construction. For example, [32, 57] finds that transforming the NER and RE task into a multi-turn question-answering dialog could improve the model performance. [9] explicitly derive syntactic knowledge to guide LLMs to think, which could develop the performance of NER. Despite these LLM-driven KG construction methods [58, 40] in general domains being widely investigated, KG construction in urban domain still remains an open challenge [59].
Urban Knowledge Graph. Urban knowledge graph has been proven useful in various urban tasks, such as traffic flow prediction [60, 61, 27, 62], mobility prediction [6], site selection [7], city profiling [63], crime prediction and so on [8, 64, 65]. Their common approach involves manually extracting urban entities and defining urban relations to construct an urban knowledge graph. For example, [6] construct a dedicated spatiotemporal knowledge graph regarding trajectory and timestamp as entities to improve trajectory prediction and [7] construct user check-in relations to help mobility prediction. Nevertheless, existing UrbanKGs heavily rely on manual design, leading to high labor costs.
7 Conclusion
In this work, we proposed UrbanKGent, the first automatic UrbanKG construction agent framework with LLMs. We first constructed a knowledgeable instruction set to adopt LLMs for different UrbanKGC tasks. Then, we proposed a tool-augmented iterative trajectory refinement module to facilitate the instruction tuning of various large language models. Extensive experimental results demonstrate the advancement of UrbanKGent in improving UrbanKGC tasks. The obtained UrbanKGent agent family, consisting of 7/8/13B version, with lower latency and cost compared with derving GPT-4 for UrbanKG construction. We hope the open-source UrbanKGent can foster future urban knowledge graph research and broader smart city applications.
Acknowledgments and Disclosure of Funding
This work was supported by the National Natural Science Foundation of China (Grant No.62102110, No.92370204), National Key R&D Program of China (Grant No.2023YFF0725004), Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.
Appendix A UrbanKGC Construction
Figure 6: Given two geo-entities X and Y, the illustrative visualization of five types of RCC relationships. In this work, we consider entities with coordinate geometry to a circles with very small radii.
Table 5: The detailed statistic of RTE datasets. We report the maximum length, minimum length, and average length of urban text in the RTE dataset.
Dataset Max Length Min Length Avg Length # Records
NYC-Instruct 1,747 68 437 232
CHI-Instruct 1,120 25 408 122
NYC 2,708 51 433 2,089
CHI 1,883 32 445 1,102
NYC-Large 4,598 20 1,179 40,480
CHI-Large 4,597 36 825 28,868
This section presents detailed UrbanKGC dataset statistic information for relational triplet extraction (RTE) and knowledge graph completion (KGC) tasks. Since two sequential tasks (i.e., RTE and KGC) of UrbanKGC are within an open-world setting (i.e., no predefined ontology) [40, 41]. Therefore, for the RTE task, every data record is an urban text without the triplet label. For the KGC task, every data record is a quadruple (i.e., head entity name, head entity geometry, tail entity name, tail entity geometry) without the geospatial relation label.
RTE Dataset. To facilitate the understanding of the constructed RTE dataset, we summarize the distribution of urban textual corpus in the six RTE datasets. As shown in Table 5, the entity distribution and textual statistics of the instruct dataset and test data are similar. In addition, the record in dataset (i.e., NYC-Instruct and CHI-Instruct) used for instruction tuning is not overlapping with that in test dataset (i.e., NYC and CHI) and the real-world application dataset (i.e., NYC-Large and CHI-Large). This can avoid potential data leakage issues.
Figure 7: The geometry range visualization of the head entity and tail entity of four KGC datasets. The horizontal and vertical coordinates are longitude and latitude, respectively. The blue and red polygons stand for entities with the polygon geometry, the purple line string stands for the entities with linestring geometry and the green point is for the coordinate entities.
KGC Dataset. As for the KGC dataset, we provide illustrative visualization of the five RCC relationships [30] in Figure 6 for better understanding. Specifically, the disconnected (DC), externally connected (EC), partially overlapping (PO), equal (EQ), tangential and non-tangential proper parts (IN) together depict the basic geospatial relationship between urban entities. Moreover, to facilitate the understanding of the constructed KGC dataset, we visualize the geometry range of the head entity and the tail entity in four small KGC datasets. Due to the large amount of data and the display overlapping between entities and entities during visualization, we will not show the visualization results of NYC-Large and CHI-Large dataset, but the pattern is similar. As can be seen in Figure 7, the entity distribution in the instruct dataset (i.e., NYC-Instruct and CHI-Instruct) and test dataset (i.e., NYC and CHI) are similar. The pattern in NYC-Large and CHI-Large is similar. Due to the KGC task follows a zero-shot setting, we are unable to provide accurate distribution proportions of the 5 RCC relationships in the four datasets. However, it is intuitive can be seen that all four datasets contain the five RCC relationships shown in Figure 6, which provides a guarantee for the practical significance of KGC task in this work.
Appendix B Instruction Template
B.1 Instruction Template in UrbanKGent
B.2 Instruction Template in Baselines
LLMs-based In-context Learning Methods. We first construct several few-shot demonstrations via the chain-of-thought prompting techniques [35], which is popular for automatic demonstration generation. The detailed instruction template is shown in Figure 9. Then, we add these demonstrations before the test question, and the detailed instruction template is shown in Figure 9.
Figure 8: The reasoning process of UrbanKGent-13B and the detailed instruction template for UrbanKent inference pipeline. The content in the gray box is the instruction, and the content in the green box is the agent’s response.
Figure 9: The instruction template of zero-shot reasoning, In-context learning and vanilla fine-tuning baselines.
Table 6: The detailed geospatial tool name, description, interface input, and output. Each tool interface is implemented by python and it is self-programmed by GPT-4.
Tool name Tool description Input Output
Geohash Geohash encoding Geometry Geohash code (8-bit)
Appendix C Geospatial Toolkit
Appendix D Evaluation
@@
This section presents the detailed evaluation process and examples to help the reader have a better understanding: (1) Human Evaluation. We employ human annotators to evaluate the results on 200 random samples. For the relational triplet extraction task, we first manually annotate the triplet label for each sample. Then, we manually evaluate the correctness of each triplet [40] based on annotation and calculate the accuracy value. For the knowledge graph completion task, we follow [11] to manually label the response as correct or wrong, and calculate the accuracy. (2) GPT Evaluation. Recently, many studies [52, 45] adopt LLM-based evaluation for open-domain tasks and empirically demonstrate that GPT-4’s evaluation and human evaluation can be consistent [66]. In this work, we also use GPT-4 to evaluate the model performance on the full data to escape intensive labor. Specifically, given an UrbanKGC instruction and results, we prompt GPT-4 to return the confidence score and the justification (i.e., True/False), which will be further used to calculate the accuracy.
Moreover, we provide a comprehensive analysis to demonstrate why GPT and Human evaluations are highly aligned.
D.1 Human Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 7. Given the urban text sentence "Columbia University is a private Ivy league research university in New York City.", the human annotators are required to first label the triplet contained in this urban text, described as triplet <Columbia University, locate-in, New York City>. Based on the label, then, they are instructed to evaluate how many true triplets and false triplets in the results from the models. Finally, they will fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). We will calculate the accuracy of results based on these annotated forms.
For the knowledge graph completion (KGC) task, given the head entity <Columbia University, Point (40.8075, 73.9626)> and the tail entity <Empire State Building, Point (40.7484, 73.9857)>, the human annotators are first required to complete their missing geospatial relationship from the five relation candidate (i.e., DC, EC, PO, EQ and IN). Specifically, the annotation could be achieved by manually visualizing the location of two entities given on the map and following the RCC relation rule in Figure 6 to determine their geographical relationships. Finally, they will fill out the evaluation results (i.e., True/False). We will calculate the accuracy based on these evaluation results.
Table 7: Illustrative RTE evaluation example when we utilize human evaluation and GPT evaluation method. We calculate the accuracy by counting the proportion of true triplets. The label for GPT evaluation method is invisible.
Type Urban text Results Label
Number of the
true triplet
Number of the
false triplet
Human
Evaluation
<Columbia University,
locate-in, New York City >
1 0
GPT
Evaluation
- 1 0
D.2 GPT Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 7. Given the urban text sentence "Columbia University is a private Ivy league research university in New York City.", and the model extraction results (i.e., <Columbia University, locate-in, New York City>), we directly prompt GPT-4 to fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). Then, the accuracy could be obtained based on these self-evaluated results.
For the knowledge graph completion (KGC) task, given the head entity <Columbia University, Point (40.8075, 73.9626)> and the tail entity <Empire State Building, Point (40.7484, 73.9857)> and completed geospatial relationship, we directly prompt GPT-4 the justify if the results are correct. Specifically, we will first explicitly call eight external geospatial tools in Table 6, and combine all these eight calculation results into the prompt. Then we feed these prompts into GPT-4 to help it make a comparable accurate evaluation. Finally, we obtain the evaluation results (i.e., True/False) for every record. We will calculate the accuracy based on these evaluation results.
D.3 Evaluation Consistency
We conduct experiments to assess the consistency between GPT-4’s evaluation results and human evaluation results. In this work, we follow GPTScore [67] to use Spearman coefficient [68] to investigate the correlation between GPT evaluation and human evaluation. Based on the evaluation results in Tabel 3, we can conduct correlation analysis separately by the types of model and paradigm.
Figure 10: The Spearman correlation between the GPT evaluation and human’s evaluation under five different LLM backbones (i.e., Llama-2-7B, Llama-2-13B, Llama-2-70B, GPT-3.5 and GPT-4).
Figure 11: The Spearman correlation between the GPT evaluation and human’s evaluation under four LLM paradigms (i.e., Zero-shot learning (ZSL), In-context learning (ICL), Vanilla fine-tuning (VFL) and UrbanKGent inference).
Appendix E In-Depth Analysis
In this section, we aim to perform more comprehensive analysis of our proposed UrbanKGent agent and hope to answer the following research questions. (1) RQ1: How does the constructed UrbanKGent perform compared with existing paradigms on larger real-world dataset? (2) RQ2: How do different components (e.g., the knowledgeable instruction generation) affect the performance? (3) RQ3: How the complexity and efficiency of proposed UrbanKGent framework? (4) RQ4: How about the UrbanKGent inference trajectories when completing UrbanKGC tasks? (5) RQ5: How can the constructed UrbanKGent provide application service to real-world scenarios?
Table 8: The average Spearman correlation value between human evaluations and GPT-4 evaluations. "Repeat X times" refers to instructing GPT-4 to generate judgments X times, and adopting the answer that appears most frequently (e.g., True/False for the KGC task and Number of the true triplet for the RTE task) as the final decision.
Task Repeat Once Repeat 3 times Repeat 5 times Repeat 10 times
RTE 98.56% 87.31% 86.42% 85.27%
KGC 94.86% 91.23% 90.09% 91.24%
E.1 RQ1: Evaluation on Larger Dataset
As shown in Table 9, we derive UrbanKGent-13B for urban knowledge graph construction using constructed large-scale dataset in NYC and CHI. Specifically, we directly use the text record in NYC-Large and CHI-Large for the relational triplet extraction task. Then, we randomly sample the head-tail entity pairs (both of head and tail entities contain geometry information) from these triplets for knowledge graph completion. Nevertheless, iterating all head-tail entity pairs is a time-consuming task, so we just construct a KGC dataset consistent with the sacle of the RTE dataset. By performing RTE and KGC tasks, we obtain two large-scale UrbanKGs shown in Table 4. Compared with existing dataset and benchmark UUKG [8], we can clearly observe that our agent can only use one-fifth data to construct the UrbanKG with the same scale entities and triplets, but extend the relationship types to a thousand times. It is worth noting that all construction process is completed by a LLM agent without any mannul effort. We think it is the core of this work.
Table 9: The experimental results of relational triplet extraction (RTE) and knowledge graph completion (KGC) on the NYC-Large dataset and CHI-Large dataset. To save the cost, We choose the best approaches in zero-shot-learning (ZSL), In-context-learning (ICL), Vanilla Fine-tuning (VFT), and UrbanKGent Inference setting.
Models NYC-Large CHI-Large
GPT(acc/confidence) Human(acc) GPT(acc/confidence) Human(acc)
RTE KGC RTE KGC RTE KGC RTE KGC
ZSL 0.36/4.01 0.45/3.95 0.42 0.31 0.38/4.27 0.33/3.67 0.44 0.36
ICL 0.39/4.36 0.48/4.15 0.48 0.39 0.38/4.22 0.35/3.53 0.40 0.35
VFT 0.37/4.05 0.46/3.98 0.45 0.35 0.33/4.25 0.28/3.67 0.38 0.34
UrbanKGent Inference
0.43/4.13 0.51/3.87 0.51 0.43 0.47/4.33 0.43/3.63 0.49 0.42
UrbanKGent-7B 0.44/4.27 0.50/4.07 0.53 0.44 0.48/4.25 0.42/3.85 0.55 0. 46
UrbanKGent-8B 0.43/4.07 0.51/4.16 0.52 0.44 0.49/4.21 0.43/3.77 0.53 0. 44
UrbanKGent-13B 0.46/4.56 0.52/3.67 0.55 0.46 0.55/4.29 0.49/3.24 0.58 0.49
E.2 RQ2: Ablation Studies
We conduct an in-depth analysis of the proposed instruction generation and tool-augmented iterative trajectory refinement module on the NYC dataset. Specifically, for the RTE and KGC task, we validate the effectiveness of each block by comparing the following variants: (1) UrbanKGent-7B removes knowledgeable instruction template in RTE and KGC task; (2) UKGent removes multi-view design in RTE task; (3) UrbanKGent-7B removes external geospatial tool invocation block; (4) UrbanKGent-7B removes iterative trajectory self-refinement. We summarize the results in Table 10, and obtain the following observations.
Table 10: Effect of different blocks.